The larger of two numbers is three more than twice the smaller and when the larger is subtracted from the smaller the result is -15.
Answer by stanbon(75887)

The larger of two numbers is three more than twice the smaller and when the larger is subtracted from the smaller the result is -15.
--------------------
Equations:
x = 2y+3
y-x = -15
----------------
Substitute for "x" and solve for "y":
y - (2y+3) = -15
-y-3 = -15
y = 12 (smaller)
---
Solve for "x":
x = 2y+3
x = 24+3 = 27 (larger)
========================
Cheers,
Stan H.
==============